OpenStudy (anonymous):

What is 3 hours and 40 minutes converted into a fraction (hours)?

OpenStudy (anonymous):

3 hours and 40/60= 3 hours and 2/3 rds

OpenStudy (anonymous):

40min = 40/60 =2/3 3h = 180min = 180/60 = 180/60 = 9/3 2/3 + 9/3 = 11/3